OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through (2, –1) and has a slope of –5.

OpenStudy (anonymous):

(y-y1)=m(x-x1) (y-(-1))=-5(x-2) (y+1)=-5x+10 y+5x=9